### v1

# Supernova-induced binary-interaction-powered supernovae: a model for SN2022jli

Abstract
We present 3D hydrodynamical modelling of supernova-induced binary-interaction-powered supernovae; a scenario proposed for the peculiar type Ic supernova SN2022jli. In this scenario, supernova ejecta of a stripped-envelope star impacts a close-by stellar companion, temporarily inflating the envelope. The expanded envelope engulfs the neutron star, causing strong mass accretion at super-Eddington rates. Feedback from the accretion powers the supernova light curve with periodic undulations. Our simulations capture key features of SN2022jli, both the overall decline and the superimposed undulations of the light curve. Based on our parameter study, we find that (i) the accretion feedback should be sufficiently geometrically confined and the eccentricity of the post-supernova binary orbit should be $0.7\lesssim e\lesssim 0.9$ to sustain a high accretion rate, and (ii) the viewing angle should be close to edge-on to match the low undulation amplitude ($\Delta L/L\sim 0.1$) of SN2022jli. Different combinations of parameters could account for other supernovae like SN2022mop, SN2009ip and SN2015ap. We also discuss possible explanations for other key features of SN2022jli such as the $\gamma$-ray detection at $\sim 200~{}\mathrm{d}$ and the rapid optical drop at $\sim 250~{}\mathrm{d}$. Finally, we speculate on the future evolution of the system and its relation to existing neutron star binaries.
Transient sources (1851) — Supernovae (1668) — Core-collapse supernovae (304)
software: HORMONE (Hirai et al., 2016; Hirai & Podsiadlowski, 2022), MESA (Paxton et al., 2011, 2013, 2015, 2018, 2019), MESA SDK (Townsend, 2021), gnuplot (Williams et al., 2021)
{CJK*}
UTF8ipxm
1 Introduction
[…]
The colour in the second portion of the light curve was much bluer than other known stripped-envelope SNe, indicating an extra powering mechanism (Cartier et al., 2024). Along with the clear periodicity, it was speculated that this SN is binary-interaction-powered as proposed in Hirai & Podsiadlowski (2022). In this scenario, the nascent neutron star (NS) directly penetrates the outer layers of a main sequence companion star, causing strong mass accretion that powers the light curve from inside the ejecta. However, such direct interactions can only occur several times as the orbit will rapidly decay upon each penetration, whereas SN2022jli showed an almost fixed periodicity for at least $\gtrsim 15$ cycles.
In this paper, we model a scenario we proposed in previous studies (Ogata et al., 2021; Moore et al., 2023; Hirai, 2024) with 3D hydrodynamic simulations. It builds on the scenario of Hirai & Podsiadlowski (2022) but with the additional effect of ejecta-companion interaction, where the outer layers of the companion star is heated by the impact of the SN ejecta and forced to inflate (e.g. Hirai et al., 2018). With sufficient SN-heating, the companion can inflate enough to overflow its Roche lobe or engulf the entire post-SN orbit, inducing binary interactions. The density of the inflated layers are typically low enough to avoid any measurable amount of orbital decay, but simultaneously high enough to enable strong mass accretion onto the remnant compact object. Since the binary orbit is almost guaranteed to be eccentric due to the SN mass loss and natal kick, we expect there to be periodic modulations in the accretion rate as the accretor travels through various depths of the companion envelope. The envelope is sufficiently optically thick to reprocess any radiative and kinetic feedback from the accretor into optical wavelengths. For all of the above reasons, this is an ideal situation to create bright undulating light curves like that of SN2022jli.
This paper is organized as follows. We first outline our model parameters and the methods for the hydrodynamic simulations and light curve modelling in Section 2. Simulation results are presented Section 3, along with a comparison of the simulated light curve with the observed light curve of SN2022jli. We then discuss how to explain other key features of SN2022jli within our scenario as well as speculations on the future evolution of the system in Section 4. We finally summarize our results in Section 5.
2 Method
2.1 Binary parameters
[…]
There are several possible constraints on the companion mass. Since the SN was classified as type Ic and the SN2022jli progenitor system likely had a pre-SN orbital period of $P_{\mathrm{orb}}<12.5~{}\mathrm{d}$, it most likely experienced a common-envelope phase that stripped the primary envelope and tightened the orbit. Although the exact conditions for initiating common-envelope phases are not known (e.g. Ivanova et al., 2013; Pavlovskii et al., 2017; Temmink et al., 2023; Willcox et al., 2023), it is considered that low mass ratio systems are generally more unstable. This places a loose upper bound on the companion mass to be sufficiently smaller than the pre-common-envelope primary mass. On the other hand, the companion mass should be large enough such that the binary can stay bound after the SN. In the absence of NS natal kicks, the binary will survive when the ejecta mass is less than half of the total mass of the binary. The estimated ejecta mass for SN2022jli is $M_{\mathrm{ej}}\sim 1.5~{}M_{\odot}$ (Chen et al., 2024; Cartier et al., 2024), so assuming a NS mass of $M_{\mathrm{NS}}=1.4~{}M_{\odot}$, we cannot place a useful lower limit on the companion mass. Given these considerations, we choose a companion mass of $M_{2}=5~{}M_{\odot}$ as our fiducial model. For comparison, we also run models with companion masses of $M_{2}=3~{}M_{\odot}$ and $10~{}M_{\odot}$. All stellar models are constructed with the stellar evolution code MESA (Paxton et al., 2011), assuming that the companion is close to its zero-age main-sequence structure.
For the orbital eccentricity, we choose $e=0.5$ as our baseline model and run $e=0.3$ and $0.7$ for comparison. All model parameters are summarized in Table 1. If we assume the pre-SN orbit was circular and that the ejecta mass was $M_{\mathrm{ej}}=1.5~{}M_{\odot}$, the mass loss alone can only excite an eccentricity of $e=M_{\mathrm{ej}}/(M_{\mathrm{NS}}+M_{2})=1/3,3/13,2/23$ for $M_{2}=3,5,10~{}M_{\odot}$ respectively. Our choices of eccentricity implies that there was also a natal kick imparted to the NS with a magnitude of the order of $v_{\mathrm{kick}}\gtrsim 50~{}\mathrm{km~{}s}^{-1}$.
Table 1: Model parameters.
Name $M_{2}$ $e$ $a_{\mathrm{per}}$ Feedback
($M_{\odot}$) ($R_{\odot}$)
M05e05b 5 0.5 21.1 Bipolar
M05e03b 5 0.3 29.6 Bipolar
M05e07b 5 0.7 12.7 Bipolar
M10e05b 10 0.5 25.6 Bipolar
M03e05b 3 0.5 18.7 Bipolar
M05e05n 5 0.5 21.1 None
M05e05t 5 0.5 21.1 Thermal
2.2 Hydrodynamics code
For all simulations we use the 3D (magneto-) hydrodynamics code HORMONE (Hirai et al., 2016), which is a grid-based code that solves the Euler equations based on a Godunov-type scheme. Self gravity is implemented via the hyperbolic self-gravity method (Hirai et al., 2016) to achieve high computational efficiency. We use an equation of state that includes the contribution of ideal gas and radiation, assuming local thermodynamic equilibrium (LTE) between the gas and radiation. HORMONE has recently been fully MPI-parallelized, enabling it to scale almost linearly on multiple compute nodes.
We use a spherical coordinate system for all of our simulations. The companion star model is centered on the origin. Since our code cannot handle vacuum, we attach a dilute atmosphere outside the star, with a density profile $\rho=\rho_{0}(R_{2}/r)^{2}$, where $\rho_{0}$ is set to a value lower than the surface density of the companion. The outer boundary extends out to $r_{\mathrm{out}}=8\times 10^{15}~{}\mathrm{cm}\sim 115,000~{}R_{\odot}$, which is $\sim 3000$ times larger than the orbital semimajor axis. The radial grid spacing is chosen such that it is smallest around the surface of the star and increases in a geometric series inwards and outwards. This ensures that the pressure scale height in the star is always resolved by at least $\sim 10$ grid points [^3] Except for the very surface layers where the pressure scale height is prohibitively small.. We cover the star with $\sim 300$ radial grid points, and $N_{r}=800$ grid points altogether. We assume equatorial symmetry, and thus only simulate the upper hemisphere where the polar grid spacing is uniform in $\cos\theta$, with $N_{\theta}=32$ grid points from pole to equator. This ensures that the solid angle per cell is uniform, minimizing asymmetries in numerical viscosity. It also places more resolution around the orbital plane, where most of the interesting dynamics are expected to take place. The azimuthal spacing is uniform with $N_{\phi}=256$ grid points. To relax the severe Courant conditions around the coordinate origin, we gradually reduce the angular resolution towards the centre, assuming full spherical symmetry ($N_{\theta}=N_{\phi}=1$) in the innermost 5 cells, $(N_{\theta},N_{\phi})=(2,16)$ in the next 5 cells, $(N_{\theta},N_{\phi})=(4,32)$ in the next 6 cells and so on.
In order to efficiently resolve the companion star at all times, we solve the hydrodynamics on a non-inertial frame fixed to the centre of the companion. To do this, at each time step, we first compute the acceleration at the coordinate origin and then subtract that value from every cell in the computational domain. We start all simulations from periastron, assuming the SN happened at $t=0$ and the orbit was instantaneously transformed to its current state.
2.3 SN-heating
After the SN explosion, the SN ejecta collide with the companion star, heating the surface layers through shocks. The injected heat drives the surface out of thermal equilibrium, causing it to become overluminous and inflated (Hirai & Yamada, 2015; Hirai et al., 2018; Ogata et al., 2021; Chen et al., 2023; Hirai, 2023). Depending on the amount of energy injection and the stellar mass, the star can reach radii of a few ×$\times 100~{}R_{\odot}$. This is far beyond the orbital semi-major axis, inevitably causing direct interactions between the NS and the inflated envelope.
In our simulations, we do not simulate the SN ejecta–companion interaction but instead inject energy into the envelope to mimic the SN-heating effect. According to the 2D hydrodynamic simulations in Hirai et al. (2018), the excess specific energy distribution takes the form
$\Delta\epsilon(m)=\frac{E_{\mathrm{heat}}}{m_{\mathrm{h}}[1+\ln{(M_{2}/m_{\mathrm{h}})]}}\times\begin{cases}1,&\text{if $m\leq m_{\mathrm{h}}$}.\\ m_{\mathrm{h}}/m,&\text{if $m>m_{\mathrm{h}}$}.\end{cases}$ (1)
where $m$ is the mass from the surface, $M_{2}$ is the companion mass, $E_{\mathrm{heat}}$ is the energy injected into the star, and $m_{\mathrm{h}}$ is a parameter that describes the efficiently heated mass. The injected energy can be computed as $E_{\mathrm{heat}}=pE_{\mathrm{exp}}\tilde{\Omega}$, where $p$ is the energy deposition efficiency, $E_{\mathrm{exp}}$ is the explosion energy and $\tilde{\Omega}$ is the fractional solid angle subtended by the companion described as
$\tilde{\Omega}\equiv\frac{1-\sqrt{1-(R_{2}/a)^{2}}}{2}\sim\frac{R_{2}^{2}}{4a^{2}}.$ (2)
[…]
The heating timescale is roughly of the order of the shock crossing time of the star
$\tau_{\mathrm{heat}}\sim 210~{}\mathrm{s}\left(\frac{R_{2}}{3\,R_{\odot}}\right)\left(\frac{v_{\mathrm{ej}}}{10^{4}~{}\mathrm{km~{}s}^{-1}}\right)^{-1},$ (3)
where $v_{\mathrm{ej}}$ is the ejecta velocity. This is much shorter than the orbital period, so we ignore this time delay and inject excess energy into the companion at $t=0$ (first periastron).
2.4 Feedback method
Accretion onto NSs is an extremely complex phenomenon involving interaction between the accreting matter and magnetic field, formation of viscous accretion disks, launching of jets and kinetic outflows, etc. Given the computational limitations, we do not attempt to resolve these phenomena, but instead treat the newly born NS as a point particle that only interacts with the gas through gravity. We soften the gravitational potential around the particle according to the cubic-spline kernel in Price & Monaghan (2007) to avoid singularities. The softening radius is chosen to be $R_{\mathrm{soft}}=1~{}R_{\odot}$ or 3 times the grid resolution around the particle, whichever is larger. The latter condition is chosen only close to apastron, where the local grid resolution is coarser, but also where the density is lower and hence the accretion is less important.
Accretion feedback is injected around the NS artificially. To assess the uncertainties of our chosen feedback form, we compare results with three types of feedback; (i) no feedback, (ii) thermal feedback and (iii) bipolar kinetic feedback. All feedback is applied in an operator split manner, where all the accretion-related terms are applied after the hydrodynamics step.
The accretion and feedback is treated in our code as follows. For all cells within a specified accretion radius $|\bm{r}-\bm{r}_{\mathrm{NS}}|\leq R_{\mathrm{acc}}$, we add a source term to the continuity equation in the form
$\frac{\partial\ln\rho}{\partial t}=-\frac{1}{t_{\mathrm{ff}}},$ (4)
where $\rho$ is the density and $t_{\mathrm{ff}}$ is the free-fall time defined by
$t_{\mathrm{ff}}\equiv\frac{R_{\mathrm{acc}}}{\sqrt{\phi_{\mathrm{soft}}(R_{\mathrm{acc}})-\phi_{\mathrm{soft}}(|\bm{r}-\bm{r}_{\mathrm{NS}}|)}},$ (5)
where $\phi_{\mathrm{soft}}(r)$ is the softened gravitational potential of the NS. Eq. (4) is based on the assumption that the mass inside the accretion radius will be accreted by the NS on the local free-fall timescale. The mass and momentum lost in the process is added to the NS particle to represent accretion. We then update the internal energy density of each cell such that the specific entropy is unchanged after applying the density source term.
For the feedback models, we further calculate the accretion luminosity as
$L_{\mathrm{acc}}=\frac{GM_{\mathrm{NS}}\dot{M}_{\mathrm{NS}}}{R_{\mathrm{NS}}},$ (6)
where $G$ is the gravitational constant, $\dot{M}_{\mathrm{NS}}$ is the mass accretion rate and we choose the NS radius to be $R_{\mathrm{NS}}=12~{}\mathrm{km}$. In the thermal feedback model, we assume that all of the accretion energy is liberated in the form of radiation, and a fraction of it is captured by the gas and thermalizes. We define the optical depth within the accretion radius as
$\tau_{\mathrm{acc}}=\kappa\bar{\rho}_{\mathrm{acc}}R_{\mathrm{acc}},$ (7)
where $\kappa$ is the opacity and $\bar{\rho}_{\mathrm{acc}}$ is the average density within the accretion radius. The energy captured by the gas is estimated as
$L_{\mathrm{therm}}=(1-e^{-\tau_{\mathrm{acc}}})L_{\mathrm{acc}}.$ (8)
This luminosity is added uniformly within the accretion radius as a source term to the energy equation. For the value of opacity, we choose $\kappa=0.1~{}\mathrm{cm}^{2}~{}\mathrm{g}^{-1}$.
In the bipolar feedback model, we assume that the accreted material forms an accretion disk that converts all of the accretion energy into bipolar outflows through jets and/or disk winds. At each time step, we first integrate the mass $M_{\mathrm{cone}}$ and total energy $E_{\mathrm{cone}}$ in the cells that are within the accretion radius ($|\bm{r}-\bm{r}_{\mathrm{NS}}|\leq R_{\mathrm{acc}}$) and satisfy
$\cos^{-1}\left(\frac{|(\bm{r}-\bm{r}_{\mathrm{NS}})\cdot\hat{\bm{n}}_{\mathrm{jet}}|}{|\bm{r}-\bm{r}_{\mathrm{NS}}|}\right)\leq\theta_{\mathrm{jet}},$ (9)
where $\hat{\bm{n}}_{\mathrm{jet}}$ is a unit vector aligned with the kinetic outflow (here we assume it is aligned with the orbital angular momentum) and $\theta_{\mathrm{jet}}$ is the opening angle of the outflow. Then for the same cells, we replace the velocity with $v_{\mathrm{jet}}(\bm{r}-\bm{r}_{\mathrm{NS}})/|\bm{r}-\bm{r}_{\mathrm{NS}}|$, where $v_{\mathrm{jet}}\equiv\sqrt{2(E_{\mathrm{cone}}+L_{\mathrm{acc}}\Delta t)/M_{\mathrm{cone}}}$ and $\Delta t$ is the time step. For the models we present, we choose $\theta_{\mathrm{jet}}=45^{\circ}$, although we tested that the results are not sensitive to the opening angle at least within the range $30^{\circ}\leq\theta_{\mathrm{jet}}\leq 45^{\circ}$.
2.5 Light curve modelling
We use the post-processing module developed for HORMONE in Grishin et al. (2021) to compute synthetic light curves. It follows the method of Suzuki et al. (2019) where we shoot a grid of rays from behind the source towards the observer and solve the radiative transfer equations to obtain the intensity along each ray. The intensity is integrated over the whole projected area of the source and multiplied by $4\pi$ to obtain the isotropic-equivalent luminosity $L$. We separately carry out the same shooting procedure for specific intensities at 6 different frequencies to get a rough spectrum. We fit the spectrum with a black body to compute an effective temperature $T_{\mathrm{eff}}$, which is then used to define a photospheric radius $R_{\mathrm{eff}}=\sqrt{L/(4\pi\sigma T_{\mathrm{eff}}^{4})}$ where $\sigma$ is the Stefan-Boltzmann constant.
For the opacity, we use an approximate analytical formula similar to Metzger & Pejcha (2017). See Appendix A for details of our opacity formula. A key component is the floor opacity, which is a commonly used method to crudely account for other unaccounted sources such as line opacities in velocity gradients, non-LTE effects, etc (e.g. Herzig et al., 1990; Bersten et al., 2011). To bracket the range of uncertainties, we compared two calculations with $\kappa_{\mathrm{floor}}=0.01$ and $0.1~{}\mathrm{cm^{2}~{}g}^{-1}$, but find that the results are mostly the same. For the rest of this paper, we show results for $\kappa_{\mathrm{floor}}=0.1~{}\mathrm{cm^{2}~{}g}^{-1}$.
Given the high asymmetry of our problem, we also investigate the viewing direction dependence. Defining a Cartesian coordinate system with the orbital angular momentum vector as the $z$ axis, the direction opposite of the orbital eccentricity vector (Laplace-Runge-Lenz vector) as the $x$ axis, and the direction of the initial NS velocity as the $y$ axis, we choose 5 different viewing directions; $+z,+x,-x,+y,-y$ (see the arrows in Figure 1). Note that the $+z$ and $-z$ directions are identical as we assume equatorial symmetry. For the ±$\pm x,\pm y$ directions, we compute the light curve from two different polar angles, $\theta=\pi/2$ (edge on) and $\pi/4$ (off plane).
We do not simulate the SN explosion itself, which would be contributing significantly to the light curve, especially at early times. We discuss implications of this in Section 4.2.
3 Results
3.1 Accretion dynamics
In Figures 13, we display snapshots of our hydrodynamical simulations. Immediately after the start of the simulation, the companion star swells up in response to the energy injection representing SN ejecta–companion interaction. The inflated envelope extends far beyond the location of the NS, creating a high-density environment around it. This leads to an extremely high mass accretion rate onto the NS, which is plotted in Figure 4. At the same time, there is almost no orbital decay from the dynamical friction between the envelope gas and NS particle due to the relatively small amount of mass in the inflated envelope ($\sim 10^{-4}$$10^{-2}~{}M_{\odot}$).
In the no feedback model (Figure 1), the inflated envelope is stirred up by the orbiting NS, creating shocks in the wake. Although part of the envelope material is sucked out of the simulation through accretion, the high density around the NS is maintained. Initially, the accretion rate onto the NS reaches as high as $\dot{M}\sim 10^{-2}~{}M_{\odot}~{}\mathrm{yr}^{-1}$, but slowly declines over time by $\sim 30~{}\%$ per orbit. Within each orbit, as the NS orbits through different depths of the envelope, the mass accretion rate (Figure 4, dark green curve) fluctuates with an amplitude of factor $\sim 5$. The accretion rate sharply rises towards each periastron and slowly declines towards apastron, loosely resembling the saw-tooth shape in the SN2022jli light curve.
Overplotted in Figure 4 is the Bondi-Hoyle-Lyttleton accretion rate computed as
$\dot{M}_{\mathrm{BHL}}=\frac{\pi G^{2}M_{\mathrm{NS}}^{2}\rho_{\infty}}{(v_{\mathrm{NS}}^{2}+c_{s,\infty}^{2})^{3/2}},$ (10)
where $v_{\mathrm{NS}}$ is the NS velocity relative to the companion star (Hoyle & Lyttleton, 1939; Bondi & Hoyle, 1944). For the density and sound speed $\rho_{\infty},c_{s,\infty}$, we use values on the opposite side of the star at the same radius as the NS to obtain representative undisturbed values of the inflated envelope. Although our method does not prescribe the Bondi-Hoyle-Lyttleton rate, our accretion rates closely follow it especially around periastron.
Figure 1: Animation of the 3D hydrodynamic simulations for the no feedback model (M05e05n). We show several different slices ($xy,xz,yz$), all cut through the centre of the companion star. The $z$ axis is the direction of the orbital angular momentum and the $x$ axis is aligned with the orbital major axis. The light blue dot marks the location of the point particle representing the NS. The animation runs from the start to end of our simulation ($\sim 100~{}\mathrm{d}$). The static version displays four select snapshots from the second orbit.
In the thermal feedback model (Figure 2), a low-density cavity is formed behind the NS. It is roughly spherical around periastron (panel (1)), but elongates into a large cone at and after apastron (panels (3)–(4)). The cavity severely quenches the accretion rate, by 4similar-toabsent4$\sim$ 4∼ 45555 orders of magnitude, down to around the Eddington limit (Figure 4, light green curve). The effect of such “negative feedback” has been discussed in depth in the context of common-envelope evolution (e.g. Soker, 2016; Grichener et al., 2021; Hillel et al., 2022).
In the bipolar feedback model (Figure 3), a low-density cavity is created in the vertical direction, while the density on the orbital plane is sustained high. The accretion rate is similarly quenched compared to the no feedback model, but only by an order of magnitude around periastron. Around apastron, the accretion is almost completely shut off similar to the thermal feedback model. As a result, the fluctuation amplitude of the mass accretion rate reaches 3similar-toabsent3$\sim$ 3∼ 36666 orders of magnitude. The peak accretion rates steadily decline for each periastron passage, at a rate similar to the no feedback model.
Figure 2: Same as Figure 1 but for the thermal feedback model (M05e05t).
Figure 3: Same as Figure 1 but for the bipolar feedback model (M05e05b).
Figure 4: Mass accretion rate evolution in our hydrodynamic simulations. We compare three different models with the same orbital parameters but different feedback methods: no feedback (model M05e05n; green), thermal feedback (model M05e05t; light green) and bipolar feedback (M05e05b; dark grey). Dashed vertical lines mark the periastron passage timings. For the no feedback model, we overplot the Bondi-Hoyle-Lyttleton accretion rate computed from Eq. (10) (dotted).
Figure 5 compares the accretion rate evolution for three models with different companion star masses. All models were computed with bipolar feedback and an orbital eccentricity $e=0.5$. They all qualitatively show the same behaviour, where there are periodic fluctuations to the accretion rate with a steady decline in the long term. There seems to be no obvious trend with mass, with the $5~{}M_{\odot}$ model showing the highest average accretion rate compared to the $3$ and $10~{}M_{\odot}$ models.
Figure 5: Same as Figure 4 but comparing models with different companion masses: $M_{2}=3~{}M_{\odot}$ (model M03e05b; sky blue), $5~{}M_{\odot}$ (model M05e05b; dark grey) and $10~{}M_{\odot}$ (model M10e05b; blue). All models were simulated with bipolar feedback.
Orbital eccentricity has a stronger influence on the accretion rate. As we can see in Figure 6, the peak accretion rates at periastron clearly correlate with eccentricity, where the $e=0.7$ model has almost an order of magnitude higher accretion rate than the $e=0.5$ model. Interestingly, the $e=0.3$ model seems to lose the clear periodicity and instead shows large fluctuations that are less obviously related to the orbital phase. Note that there is a two-fold effect of varying eccentricity in our setup. Since we fix the orbital period to $P_{\mathrm{orb}}=12.5~{}\mathrm{d}$, the eccentricity determines the periastron distance and hence the pre-SN orbital separation. Higher eccentricity leads to more energy injection through ejecta-companion interaction (for a given explosion energy), as the solid angle subtended by the companion is larger. Therefore, the higher eccentricity models lead to larger inflated radii with more mass in the inflated envelope. The the other effect of higher eccentricity is that with shorter periastron distances, NS will orbit in deeper parts of the envelope. Even if the envelope profiles were the same, the higher eccentricity models would have higher mass accretion rates because the NS reaches down to the deeper, higher density layers.
Figure 6: Same as Figure 4 but comparing models with different eccentricities: $e=0.3$ (model M05e03b; pink), $0.5$ (model M05e05b; dark grey) and $0.7$ (model M05e07b; red). All models were simulated with bipolar feedback.
We also compare the mass ejected from the companion star in Figure 7. To compute the companion mass, we integrate the mass in all bound cells, where a cell is marked as bound when $\phi_{\mathrm{gas}}+v^{2}/2<0$. Here, $\phi_{\mathrm{gas}}$ is the gravitational potential of the gas that does not include the contribution from the NS particle, and $v\equiv|\bm{v}|$ is the cell velocity in the frame of the companion. We see that straight after the SN, $10^{-3}$$10^{-2}~{}M_{\odot}$ of mass is ejected from the companion due to the SN-heating. While only $2\times 10^{-3}~{}M_{\odot}$ is ejected in the $e=0.3$ model (pink curve), $1.2\times 10^{-2}~{}M_{\odot}$ is ejected in the $e=0.7$ model (red curve) due to the larger energy injection via ejecta-companion interaction. The curves with the same orbital parameters but different feedback methods (grey, dark green and light green curves) show almost identical initial mass loss (Figure 4). This confirms that the initial jump is purely due to SN-heating and not because of subsequent energy injection from the feedback. After the second periastron passage, the ejected mass show small deviations between the feedback and no-feedback model, indicating that accretion feedback is causing more mass ejection from the system. Given that the total mass accretion over the duration of our simulations is $\Delta M_{\mathrm{NS}}\lesssim 10^{-4}~{}M_{\odot}$, we can safely say that this mass is ejected from the system and not accreted onto the NS. In the $e=0.7$ model, we can see more significant mass ejection episodes around each periastron passage. The mass ejected per periastron decreases each orbit, with $\sim 10^{-3}~{}M_{\odot}$ ejected in the first passage while only a few $\sim 10^{-4}~{}M_{\odot}$ is ejected in the later orbits.
Figure 7: Time evolution of the mass ejected from the donor star. We compare all models with $M_{2}=5~{}M_{\odot}$. We treat model M05e05b as the “Baseline” model and colour variations are the same as in Figures 4 and 6.
3.2 Light curve
Figure 8 displays our synthetic light curves based on the hydrodynamic simulations. We show the M05e07b model as it has the highest accretion rate among our models with feedback, but all other models show qualitatively similar behaviour. The grey curve shows the accretion luminosity from the NS particle, which is directly proportional to the mass accretion rate (Eq. (8)). Due to the strong phase-dependence of the accretion, the energy injection into the system can be considered as an almost impulsive explosion at each periastron passage. The resulting light curves have large variations depending on the viewing direction, reflecting the complex geometry of the source binary.
Right after the SN explosion, there is almost no viewing direction dependence. This is because initially, the companion envelope that reprocesses the injected energy expands in an almost spherical manner due to our SN-heating method. After the second periastron passage, the light curves start to significantly deviate from each other due to the envelope being distorted by the orbital interaction. During the first few orbits, the peak luminosities are highest in the $-x$ (pink curve) and $+y$ (blue curve) directions. These angles respectively view the binary from the equatorial plane in the direction of periastron ($-x$) and the direction perpendicular to that on the post-periastron side ($+y$). From this side of the orbit, the NS is in front of the companion star when the accretion rate (and therefore the feedback) is high. On the other side of the orbit ($+x$ and $-y$ directions), the NS is behind the companion when the accretion rate is high, so the radiation is damped by the optically thick inflated envelope before it reaches the observer.
Each successive peak generally gets dimmer and dimmer for two reasons. One is that the accretion rate declines every periastron due to the decline in envelope density. Another effect is the build-up of equatorial obscuring material due to multiple mass ejections at periastron passage. Such mass ejections only occur along the orbital plane, and conversely in the poleward direction (±$\pm z$), the material gets more and more evacuated by the bipolar outflows. This leads to a rise in the peak luminosities as we can see in the purple curve.
Figure 8: Simulated light curves for the $e=0.7$ model (M05e07b). Each coloured curve shows the isotropic-equivalent bolometric luminosity from various viewing directions as defined in Figure 1. Dashed curves show the light curve from intermediate viewing angles $\theta=\pi/4$ ($45^{\circ}$ off the orbital plane) of the corresponding directions. The grey curve shows the accretion luminosity from the NS particle (Eq. (8)) and the black crosses are the observed light curve of SN2022jli (Chen et al., 2024). The brown shaded region covers the phase where we expect the main SN ejecta to be optically thick and thus our accretion-powered light curve is still not visible (see Section 4.2).
Figure 9: (Upper panel) Effective temperature from our light curve simulations for various viewing directions. (Lower panel) Photospheric radius computed from the simulations. Colours of the curves indicate the same viewing directions as in Figure 8.
All light curves show some level of undulation, where the amplitude strongly depends on the viewing direction. Generally speaking, both the peak luminosity and undulation amplitude is smaller when viewed from the orbital plane compared to the poles. When viewed from intermediate viewing angles (dashed curves), the light curve is generally in between the pole-on light curve (±$\pm z$) and the equatorial light curve of the given azimuthal angle $\varphi$.
Overall, our light curves resemble the key features of the SN2022jli light curve. After $\gtrsim 50~{}\mathrm{d}$, the light curve reaches a state where it steadily declines with some periodic undulations. This is likely the most important characteristic of SN2022jli that sets it apart from other known transients. Typical SNe show declines but without periodic undulations. Some mass-transferring binaries, such as X-ray binaries, show periodic modulations without a long-term decline. The combination of both the decline and undulation is naturally achieved in our model. Quantitatively, the bolometric luminosity is in the same ballpark as that of SN2022jli and so is the decline rate. The undulation amplitude is also consistent if viewed from close to the orbital plane. On the other hand, there are some key features that we do not reproduce, including the first $\lesssim 50~{}\mathrm{d}$ and the drop at $\sim 250~{}\mathrm{d}$. We discuss these points in Section 4.
3.3 Photospheric evolution
As we show in Figure 9, the effective temperature steadily decreases while the photospheric radius gradually increases over time in our models. There are periodic signatures both in the temperature and radius, which become weaker for the edge-on viewing directions. The rise in photospheric radius seems to stall around the fifth periastron, and may be starting to decline towards the end of our simulation. This plateau-like behaviour is due to the opacity having a steep drop around the recombination temperature $\sim 10^{4}~{}\mathrm{K}$. Both temperature and radius roughly have the same order of magnitude as observed in SN2022jli ($T_{\mathrm{eff}}\sim 10^{4}~{}\mathrm{K},R_{\mathrm{eff}}\sim 10^{15}~{}\mathrm{cm}$). The photospheric radius is still well contained within our computational domain ($8\times 10^{15}~{}\mathrm{cm}$), validating our choice of the outer boundary radius. While we do not see the overall blue-ward evolution seen in SN2022jli, the model reproduces the correct phase evolution where the colour is bluer at undulation peaks (Cartier et al., 2024).
The complex structure of the photosphere can be viewed more clearly in Figures 10 & 11. Here, we plot 1D profiles along various radial rays at given snapshots. The curves are colour-coded based on the latitudinal angle, with darker colours towards the pole and lighter colours towards the equator. We can clearly see the surface layers of the star have been significantly inflated from the initial profile (grey dashed curve), extending out to the location of the NS and beyond.
At periastron (Figure 10), we can see the low-$\theta$ curves have much lower density compared to the high-$\theta$ curves around the NS (vertical line), which is because of the bipolar cavity blown out by the feedback. The profile interior to the location of the NS is roughly spherical with low velocities, indicating it is roughly in hydrostatic equilibrium. The exterior is extremely asymmetrical, showing $\lesssim 5$ orders of magnitude variations in density depending on the direction. A similar trend can be seen at apastron (Figure 11), with now a larger hydrostatic interior region and strongly perturbed exterior. Generally, the equatorial profiles seem to have higher densities compared to the polar profiles, showing that the outflow is equatorially concentrated.
The opacity structure is also quite complex but broadly speaking it is almost like a step function, sharply transitioning from the electron scattering opacity to the floor opacity at the recombination front ($\sim 10^{4}~{}R_{\odot}$). The radius of the recombination front, which corresponds to the photosphere, varies by a factor $\sim 5$ depending on the direction. We note that the opacity jumps at the photosphere are only covered with 3–4 grid points so it is not well resolved. Together with the crude approximation for the opacity that ignores effects of lines, non-equilibrium ionization states, multi-dimensional effects, etc, the temperature and photospheric radius evolution in our models should be approached with caution.
Figure 10: Profiles of density, pressure, temperature, opacity and radial velocity as a function of distance from the companion core. Each curve corresponds to the density profile along a specific radial ray and are coloured according to the latitudinal angle. The brown vertical lines mark the location of the NS. We choose a snapshot from model M05e07b at $t=37.5~{}\mathrm{d}$, which is the third periastron. The grey dashed curves show the initial profiles.
Figure 11: Same as Figure 10 but for $t=43.75~{}\mathrm{d}$; the fourth apastron.
3.4 Undulation features
We now compare the shape of the undulations between our model and observations. We focus on the light curve for the $+x$ viewing direction in Figure 8 (red curve) as it has the smallest undulation amplitude. To extract the undulations, we first fit a sixth-order polynomial curve to the light curve between 50 and 100 d to obtain a baseline. We plot the difference between the light curve and the baseline against the undulation phase in Figure 12. The observed undulation light curve is computed in the same way and overplotted. Overall, the model is in good agreement with the observed shape. There is a sharp rise followed by a longer decline. Each cycle has slightly different shapes both in the model and observations, partly due to the ambiguity in defining the baseline, but also due to the stochastic nature of the undulations.
Figure 12: Undulation light curve folded over undulation phase. Grey points are the observed data of SN2022jli (Chen et al., 2024). Red solid curves are the folded light curves of the $+x$ model in Figure 8. Both are shifted so that the minima are located at the phase origin.
While our edge-on light curve models closely agree with the small amplitude undulations in SN2022jli, we also predict large amplitude oscillations if a similar event was viewed from higher-latitude viewing directions. Recently, another rare kind of stripped-envelope SN was discovered which re-emerged as a type IIn SN, $\sim 3$ yr after the first event (SN2022mop; Brennan et al., 2025). In both the tail of the first event and the rise towards the second event, several periodic outbursts ($\sim 1$ mag) were detected with $\sim 27~{}\mathrm{d}$ intervals. The shape and amplitude of the outbursts loosely resemble the saw-tooth like morphology that we find in our hydrodynamic models observed from the pole (purple curves in Fig 8). It may be that the first event of SN2022mop was an event similar to SN2022jli, but with different orbital periods and viewed from different angles. Brennan et al. (2025) report that SN2009ip also has periodic outbursts with amplitudes and intervals similar to SN2022mop.
4 Discussion
4.1 Super-Eddington accretion
[…]
At such high accretion rates, neutrino emission may start to dominate the cooling in the inner parts of the flow, enabling hypercritical accretion (e.g. Chevalier, 1993, 1996; Fryer et al., 1996). The accretion rate range for neutrino-dominated hypercritical steady accretion onto NSs is thought to be around $10^{-4}~{}M_{\odot}~{}\mathrm{yr}^{-1}\lesssim\dot{M}\lesssim 0.1~{}M_{\odot}~{}\mathrm{yr}^{-1}$, consistent with the accretion rate required to power the luminosity of SN2022jli. In this scenario, the accretion flow is capable of forming a neutrino-cooled disk, which could power bipolar outflows, e.g. via the Blandford-Payne mechanism (Blandford & Payne, 1982), necessary to avoid excessive negative feedback.
4.2 $\sim$50 d delay to the second peak
In our synthetic light curves, the accretion feedback occurs straight after the SN as the timescale of the companion inflation is on the dynamical timescale. As a result, the accretion feedback kicks in almost immediately, which contradicts with the $\sim 50~{}\mathrm{d}$ delay to the second peak in SN2022jli. Previous studies have shown that the first part of the light curve is consistent with being powered by ${}^{56}\mathrm{Ni}$ decay and can be modelled as a typical type Ic SN. In our hydrodynamic simulations, we have not included the contribution of the SN ejecta, which will greatly contribute to the light at early times but also could be optically thick enough to hide the radiation from the accretion feedback. The $\sim 50~{}\mathrm{d}$ delay to the second peak could therefore correspond to the time it takes for the SN ejecta to become optically thin and expose the interior binary.
The optical depth of SN ejecta can be roughly estimated as
[…]
where $\kappa$ is the ejecta opacity and $M_{\mathrm{ej}},R_{\mathrm{ej}},v_{\mathrm{ej}}$ are the ejecta mass, radius and velocity, respectively. The factor $f$ is a parameter that depends on the density distribution along the line of sight. Given that we see no signatures of circumstellar matter interaction in SN2022jli, we can safely assume that the ejecta follow a free expanison ($R_{\mathrm{ej}}=v_{\mathrm{ej}}t$). The ejecta become optically thin when $\tau_{\mathrm{SN}}\lesssim 1$, at around
$t_{\mathrm{nebular}}\gtrsim 50~{}$ $\mathrm{d}\left(\frac{f}{0.1}\right)^{\frac{1}{2}}\left(\frac{\bar{\kappa}}{0.1~{}\mathrm{g~{}cm^{-2}}}\right)^{\frac{1}{2}}$
×$\times\left(\frac{M_{\mathrm{ej}}}{M_{\odot}}\right)^{\frac{1}{2}}\left(\frac{v_{\mathrm{ej}}}{10^{4}~{}\mathrm{km~{}s^{-1}}}\right)^{-1},$ (12)
which is consistent with the $\sim 50~{}\mathrm{d}$ delay at the order of magnitude level. Indeed, typical stripped-envelope SNe reach the nebular phase within $\sim 100~{}\mathrm{d}$, roughly consistent with this estimate.
4.3 Rapid drop at $\sim$250 d
There are several possible explanations for the rapid drop at $\sim 250~{}\mathrm{d}$. First, it could be due to rapid formation of dust, which enshrouds the emitting region. Evidence of dust formation has been reported around the time of the drop (Cartier et al., 2024). Given that the SN ejecta are very C-rich (it is a type Ic SN), it is a perfect site for dust formation. As soon as the ejecta adiabatically cool below a threshold temperature, CO molecules can form, which further enhances the cooling through molecular lines. It then quickly reaches the dust condensation temperature to form dust and enshroud the inner binary. In this case, we may expect that there is continuous energy injection from the embedded binary and the radiation is reprocessed into infrared bands. Although in SN2022jli the infrared flux does increase after the optical drop ($\sim 400~{}\mathrm{d}$), it does not make up for the decrease in optical flux, indicating a drop in the energy injection is also required.
Another possibility within our scenario is that the envelope inflation ceases at some point, preventing any further mass transfer to the neutron star. The strong accretion onto the NS quickly drains the material in the inflated envelope. At some point the SN-heated envelope material will be accreted or ejected from the system, bringing an end to the accretion process. Unfortunately due to limitations in computational resources, we were only able to run the simulations up to $\lesssim 100~{}\mathrm{d}$. It will be interesting in the future to extend the simulation to see if such a drop naturally occurs.
We can also consider the so-called propeller effect. The new-born neutron star is likely rapidly spinning and possesses a magnetic field. As the accretion rate onto the neutron star declines over time, there will come a point where the Alfvén radius exceeds the corotation radius, beyond which the magnetic field will centrifugally blow away the accreting material. The critical luminosity below which the propeller mechanism kicks in can be estimated by (e.g. Campana et al., 2002)
$L_{\mathrm{lim}}\sim 3.6\times 10^{41}~{}$ $\mathrm{erg~{}s^{-1}}\xi^{\frac{7}{2}}\left(\frac{B}{10^{12}~{}\mathrm{G}}\right)^{2}\left(\frac{P_{\mathrm{spin}}}{20~{}\mathrm{ms}}\right)^{-\frac{7}{3}}$
×$\times\left(\frac{M_{\mathrm{NS}}}{1.4~{}M_{\odot}}\right)^{-\frac{2}{3}}\left(\frac{R_{\mathrm{NS}}}{10~{}\mathrm{km}}\right)^{5},$ (13)
where $B$ is the strength of the magnetic field, $P_{\mathrm{spin}}$ the spin period and $\xi$ is an order unity parameter that accounts for geometrical effects. In SN2022jli, the drop at $\sim 250~{}\mathrm{d}$ occurred when the luminosity declined below $\lesssim 3\times 10^{41}~{}\mathrm{erg}$, which corresponds to the limiting luminosity of a NS rotating at $P_{\mathrm{spin}}\sim 20$ ms for a typical magnetic field strength $B\sim 10^{12}$ G. This spin period is within the range of birth spins of NSs predicted from core-collapse simulations (e.g. Janka et al., 2022; Burrows et al., 2024). As soon as the accretion luminosity declines below the critical luminosity ($L_{\mathrm{acc}}<L_{\mathrm{lim}}$), the accretion will be magnetically inhibited, leading to a rapid drop in the accretion rate.
A final possibility is that the infall rate onto the NS falls below the threshold for neutrino-dominated accretion. It is unclear what should happen in the transition from a neutrino-dominated to radiative accretion flow, but it is possible that the nature of the accretion changes significantly. For example, as the neutrino cooling weakens, the accretion disk becomes geometrically thicker, making the jet collimation mechanism less efficient. As we see in our simulations, more spherical feedback causes stronger negative feedback, severely quenching the accretion rate.
It is possible that multiple mechanisms are responsible for the drop, as the above models are not mutually exclusive. For example, when the accretion luminosity drops, the temperature around the photosphere will also decline. This could trigger dust formation, further accelerating the drop in optical flux. As we see in later sections, the dust formation is likely a secondary effect and not the sole cause for the drop.
4.4 Gamma-ray detection
The origin of the GeV $\gamma$-rays are still unknown. Cartier et al. (2024) propose that the $\gamma$-rays could be generated from a magnetar wind nebula. In their model, the nascent NS possesses a strong magnetic field ($B\sim 8.5\times 10^{14}~{}\mathrm{G}$) as well as rapid spin ($P_{\mathrm{spin}}\sim 50~{}\mathrm{ms}$), and both the optical light curve and $\gamma$-rays are powered by the spin-down energy.
In our accretion-powered scenario, the $\gamma$-rays could be generated by a relativistic jet launched from the NS (Figure 13). High-energy ($\sim$100 MeV–TeV) $\gamma$-rays have been observed in the past in stellar mass binaries like LS I$+61^{\circ}303$ (Abdo et al., 2009; Albert et al., 2009). The compact object in this system is most likely a NS (Weng et al., 2022), indicating that NSs are capable of creating GeV $\gamma$-ray emission. While magnetar-like behaviour has been reported (Torres et al., 2012; Weng et al., 2022), bipolar jets have also been detected in radio (Massi et al., 2004, 2012), complicating the identification of the $\gamma$-ray source in LS I$+61^{\circ}303$.
Both mechanisms generate $\gamma$-rays at the shock between the NS outflow (magnetar wind/accretion-powered jet) and the surrounding material, which in this case is the inflated envelope of the companion. The optical depth of the envelope should decline under $\tau\lesssim 1$ for the $\gamma$-rays to emerge. For GeV photons, the main source of opacity is Compton scattering in the Klein-Nishina regime. The Klein-Nishina cross section is approximately $\sigma_{\mathrm{KN}}\approx 3\sigma_{\mathrm{T}}m_{e}c^{2}/(8\epsilon_{\gamma})$ (Lang, 1980; Barkov & Komissarov, 2011), where $\sigma_{\mathrm{T}}$ is the Thomson scattering cross section, $m_{e}$ is the electron mass, $c$ the speed of light and $\epsilon_{\gamma}$ the photon energy. Following Eq. (12), we can estimate the $\gamma$-ray emergence time due to Klein-Nishina scattering as
$t_{\gamma,e}\gtrsim 1.3~{}$ $\mathrm{d}\left(\frac{f}{0.1}\right)^{\frac{1}{2}}\left(\frac{\epsilon_{\gamma}}{1~{}\mathrm{GeV}}\right)^{-\frac{1}{2}}$
×$\times\left(\frac{M_{\mathrm{ej}}}{M_{\odot}}\right)^{\frac{1}{2}}\left(\frac{v_{\mathrm{ej}}}{10^{4}~{}\mathrm{km~{}s^{-1}}}\right)^{-1}.$ (14)
The ejecta would thus become transparent to GeV photons within days.
The shocks responsible for generating $\gamma$-rays could also create X-ray emission via e.g. synchrotron radiation. These photons could act as a source of opacity through photon-photon absorption, in which two photons interact to create electron-positron pairs ($\gamma+\gamma\rightarrow e^{-}+e^{-}$). To absorb GeV $\gamma$-ray photons, the target soft photons should have an energy $\epsilon_{\mathrm{soft}}\geq(2m_{e}c^{2})^{2}/\epsilon_{\gamma}\sim 1~{}\mathrm{keV}$. Although such keV photons will eventually be re-processed into optical wavelengths via Thomson scattering, they could still interact with the $\gamma$-rays before being diminished. The optical depth for photon-photon absorption can be estimated as (Barkov & Komissarov, 2011)
$\tau_{\gamma\gamma}\sim\frac{\sigma_{T}}{5}\frac{L_{\mathrm{soft}}}{4\pi R_{\mathrm{sh}}c\epsilon_{\mathrm{soft}}},$ (15)
where $L_{\mathrm{soft}}$ is the X-ray luminosity, $R_{\mathrm{sh}}$ is the radius of the $\gamma$-ray emitting shock front and $\epsilon_{\mathrm{soft}}=1~{}\mathrm{keV}$ is the X-ray energy. In this scenario, GeV emission will appear when $R_{\mathrm{sh}}$ exceeds the “absorption radius”
$R_{\mathrm{abs}}\sim\frac{\sigma_{T}}{5}\frac{L_{\mathrm{soft}}}{4\pi c\epsilon_{\mathrm{soft}}}\approx 3000~{}R_{\odot}\left(\frac{L_{\mathrm{soft}}}{10^{42}~{}\mathrm{erg~{}s}^{-1}}\right),$ (16)
below which the GeV photons will be absorbed by the keV photons. In our simulations, the polar shock front lies roughly around $R_{\mathrm{sh}}\sim 2000~{}R_{\odot}$ at $\sim 40~{}\mathrm{d}$ (see the outer edge of the polar $v\sim 10^{10}~{}\mathrm{cm~{}s}^{-1}$ component in Figures 10 & 11) and advances linearly in time. If we assume that the X-ray luminosity at the shock is equal to the observed optical luminosity, the absorption radius declines over time. The crossover of the increasing $R_{\mathrm{sh}}$ and decreasing $R_{\mathrm{abs}}$ occurs in a few months, not too far from the $\sim 200~{}\mathrm{d}$ delay in SN2022jli. See Figure 13 for a schematic diagram of the $\gamma$-ray emission in our model.
There are no reported detections of $\gamma$-rays after the rapid optical drop. If accretion is ongoing, we would expect the $\gamma$-rays to still be observable after the drop because dust is mostly transparent to $\gamma$-rays. The lack of ongoing $\gamma$-ray emission indicates that dust obscuration cannot be the sole reason for the optical drop but the accretion has likely terminated or significantly weakened around the drop.
Figure 13: Schematic diagram of the emission from SN2022jli at $\sim 200~{}\mathrm{d}$.
[…]
4.6 Orbital eccentricity
Within our scenario, we can further narrow down the possible range of eccentricity based on orbital stability. The undulation period in SN2022jli is preserved for at least $\gtrsim 200~{}\mathrm{d}$, indicating that there is no orbital decay through drag or mass transfer processes. Figure 14 displays the maximum eccentricity to avoid orbital decay, computed as $e_{\mathrm{max}}=1-R_{2}/(fa)$. For the red curves we set $f=1$ ($a_{\mathrm{per}}=a(1-e_{\mathrm{max}})=R_{2}$) and for the black curves we use the Roche lobe radius formula (Eggleton, 1983)
$f=\frac{0.49q^{2/3}}{0.6q^{2/3}+\log(1+q^{1/3})},$ (17)
where $q=M_{2}/M_{\mathrm{NS}}$ is the mass ratio. Any eccentricity in the grey shaded region would cause Roche lobe overflow at periastron, likely leading to noticeable changes in the orbital period. The red region is a harder limit where the NS will plunge into/through the companion star core, which should drive rapid orbital decay (Hirai & Podsiadlowski, 2022). Given that there is no evidence of orbital decay all the way to the drop, the system eccentricity should have been below the grey shaded region ($e\lesssim 0.88$).
Figure 14: Allowed eccentricities as a function of companion mass. The neutron star mass is fixed to $M_{\mathrm{NS}}=1.4~{}M_{\odot}$ and the orbital period to $P_{\mathrm{orb}}=12.5~{}\mathrm{d}$. Black curves show the minimum eccentricity above which the companion star will overflow its effective Roche lobe at periastron whereas the red curves show the minimum eccentricity above which the NS will plunge into the companion at periastron. We use the Hurley et al. (2000) formulae to estimate the stellar radius $R_{2}$ at zero-age main sequence (dashed) and at $t=10~{}\mathrm{Myr}$ (solid).
4.7 Future evolution
The inflation of SN-heated companion stars only last for several years to decades (Hirai & Yamada, 2015; Hirai et al., 2018; Ogata et al., 2021; Hirai, 2023). Once the star regains thermal equilibrium, it should return to a regular main-sequence star which does not overflow its Roche lobe. Our hydrodynamic simulations unfortunately cannot place strong constraints on the possible companion mass of SN2022jli (Figure 5). Here, we speculate on the future of SN2022jli depending on what the companion star mass is.
If the companion star is an OB-type star ($M_{2}\gtrsim 8~{}M_{\odot}$), the system may immediately transition into an eccentric wind-fed high-mass X-ray binary similar to GX 301-2 (Sato et al., 1986). Given that there was a likely mass transfer phase preceding the SN that stripped the hydrogen and helium layers, the companion star could have accreted a substantial amount of mass and angular momentum. Therefore, the companion may eventually evolve into a rapidly-rotating Oe/Be star, turning the system into a Be X-ray binary like LS I$+61^{\circ}303$. This system shares many properties with the SN2022jli system, such as a similar orbital period ($P_{\mathrm{orb}}\sim 26.5~{}\mathrm{d}$; Taylor & Gregory, 1982), eccentricity ($e\sim 0.54$; Aragona et al., 2009) and high-energy $\gamma$-ray emission (Abdo et al., 2009; Albert et al., 2009). In this case, SN2022jli was marking the birth of a high-mass X-ray binary, with an enhanced accretion phase straight after the SN.
[…]
5 Summary and conclusion
We present 3D hydrodynamic simulations of SN-induced binary-interaction-powered SNe and predict their observational features. We find that our scenario can robustly produce light curves with periodic undulations and a steady decline in the long term. The luminosity ($\sim$ mass accretion rate) depends strongly on various factors, such as the form of accretion feedback and orbital eccentricity.
Based on our parameter study, a high-eccentricity ($e\gtrsim 0.7$) orbit with geometrically confined feedback is required to reach mass accretion rates compatible with the luminosity of SN2022jli. In addition, the viewing angle should be close to edge-on in order to match the relatively small undulation amplitudes of SN2022jli. Higher viewing angle models are reminiscent of more strongly outbursting SNe like SN2022mop or SN2009ip. On the other hand, the companion mass does not seem to influence the accretion rate significantly.
Binary-interaction-powered SNe provide strong evidence that binary interactions like mass transfer or common envelope phases are responsible for stripping the hydrogen-rich envelope and explode as stripped-envelope supernovae (Podsiadlowski et al., 1992). By comparing our models to the observed properties of the light curve, we can further constrain the immediate post-SN binary properties. Similar to how post-SN companion detections can help constrain binary interaction physics (Hirai, 2023), the post-SN binary properties can also indirectly constrain binary interaction physics. For example, in our model for SN2022jli, the post-SN periastron distance must be $a_{\mathrm{per}}\lesssim 19~{}R_{\odot}$. This directly relates to the pre-SN orbital separation, which in turn relates to the post-binary interaction orbital separation. If the system had undergone a common-envelope phase, our constraint can provide useful insight into the long-sought post-common envelope separations in massive binary systems.
Since the first discovery of periodic undulations in a SN (SN2022jli), a growing number of possible periodic outbursts have been reported in the past few years, e.g. SN2022mop, SN2009ip (Brennan et al., 2025), SN2015ap (Ragosta et al., 2025). These all may be marking the birth of compact object binaries including X-ray binaries, detached NS binaries, etc. Many more such SNe are expected to be discovered in the coming years with next-generation transient survey telescopes including the Vera Rubin Observatory and the Wide Field Survey Telescope. In particular, the high cadence of the Legacy Survey of Space and Time (LSST) by the Rubin Observatory (especially in the deep drilling fields) could be extremely powerful for identifying periodic signatures in SN light curves. Our models should serve as a first guide into what to expect in the future, and how to infer the underlying binary properties from the observed light curves.
This work was supported by software support resources awarded under the Astronomy Data and Computing Services (ADACS) Merit Allocation Program. ADACS is funded from the Astronomy National Collaborative Research Infrastructure Strategy (NCRIS) allocation provided by the Australian Government and managed by Astronomy Australia Limited (AAL). This work was partially performed on the OzSTAR national facility at Swinburne University of Technology. The OzSTAR program receives funding in part from the Astronomy NCRIS allocation provided by the Australian Government. Part of the simulations were performed on the supercomputer Gadi with the assistance of resources from the National Computational Infrastructure (NCI Australia), an NCRIS enabled capability supported by the Australian Government. Part of the simulations were also performed on the HOKUSAI computing facility in RIKEN. This work was initiated in part at Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452. S.N. is supported by JSPS Grant-in-Aid Scientific Research (KAKENHI) (A), Grant Number JP25H00675, and by JST ASPIRE Program “RIKEN-Berkeley mathematical quantum science initiative”.
Appendix A Approximate opacity formula
For the opacity in our light curve calculations, we use an approximate analytical formula similar to that of Metzger & Pejcha (2017). We combine various approximate formulae as
$\kappa=\max\left[\kappa_{\mathrm{floor}},\kappa_{\mathrm{mol}}+\left(\kappa_{\mathrm{H}^{-}}^{-1}+\left(\kappa_{\mathrm{es}}+\kappa_{\mathrm{K}}\right)^{-1}\right)^{-1}\right],$ (A1)
where the electron scattering opacity is
$\kappa_{\mathrm{es}}=$ $~{}0.2(1+X)\left(1+2.7\times 10^{11}\frac{\rho/\mathrm{(g~{}cm^{-3})}}{(T/\mathrm{K})^{2}}\right)^{-1}$
×$\times\left(1+\left(\frac{T}{4.5\times 10^{8}~{}\mathrm{K}}\right)^{0.86}\right)^{-1}~{}\mathrm{cm^{2}~{}g^{-1}},$ (A2)
Kramers opacity (bound-free/free-free absorption) is expressed as
$\kappa_{\mathrm{K}}\approx 4\times 10^{25}Z(1+X)\left(\frac{\rho}{\mathrm{g~{}cm^{-3}}}\right)\left(\frac{T}{\mathrm{K}}\right)^{-7/2}~{}\mathrm{cm^{2}~{}g}^{-1},$ (A3)
we approximate the opacity of negative hydrogen as
$\kappa_{\mathrm{H}^{-}}\approx 1.1\times 10^{-40}Z^{0.5}\left(\frac{\rho}{\mathrm{g~{}cm^{-3}}}\right)^{0.2}\left(\frac{T}{\mathrm{K}}\right)^{11}~{}\mathrm{cm^{2}~{}g^{-1}},$ (A4)
and $\kappa_{\mathrm{mol}}\approx 0.1Z~{}\mathrm{cm^{2}~{}g^{-1}}$ is the molecular opacity. The form of $\kappa_{\mathrm{H^{-}}}$ was chosen by a fit-by-eye approach to the opacity table presented in Farag et al. (2024). We do not include dust opacities, as dust formation was only detected after $\gtrsim 250~{}\mathrm{d}$ (Cartier et al., 2024). A floor opacity $\kappa_{\mathrm{floor}}$ is set to account for various other effects including lines, velocity gradients, non-LTE effects, etc.
The resulting opacities are displayed in Figure 15. We find that for the relevant density–temperature combinations in our simulation, the opacity is dominated by electron scattering.
Figure 15: Opacity as a function of density and temperature computed from Eq. (A1) with $\kappa_{\mathrm{floor}}=0$. Overplotted are the density and temperature of each cell in our simulations from the two snapshots used in Figures 10 & 11.